# A Representation Learning Approach to Feature Drift Detection in Wireless Networks

### v1

@@ VI-E Performance Metrics
$Score=\begin{cases}F1_{gain},&\text{if decision is true positive}.\\ T_{s},&\text{if decision is true negative}.\\ F1_{gain}-T_{s},&\text{if decision is false positive}.\\ -F1_{gain},&\text{if decision is false negative}.\\ \end{cases}$ (10)
The function is designed to increase the score for correct data drift predictions and to decrease for incorrect ones. The larger the data drift, the bigger score is allocated for correct assessments, and the bigger the penalty is given for incorrect predictions. We assume that each method under examination outputs a prediction on whether the $M_{0}$ should be retrained. We distinguish $4$ different possible scenarios for this assessment:
@@
  • A method’s assessment is false negative and incorrectly identifies the absence of feature drift. The penalty of this error is $-F1_{gain}$, which is the $F1$-score that the model would gain if it was retrained with the new data.
@@ VI-F Utility Component Contribution
[…]
@@ VII-A Parameter Search for the Representation Learning Components
Figure 6 illustrates the results of the $3$-dimensional parameter exploration, which is conducted for the MLP. We experiment with two datasets in which the data drift phenomenon is prominent, in order to calibrate the MLP parameters. We should note that the utility function should be represented by a high value, to clearly predict the existence of data drifts. Results indicate that the increase of model complexity is directly correlated with higher utility scores. This is expected, since more complex models manage to properly capture input data representations and thus, to produce better assessments. Nonetheless, there is an optimal parameter space, over which larger and more complex MLP s do not provide significant contributions to the utility score. Empirically, an MLP with $3$ layers, $20$ neurons per layer, and a training period of $3$ epochs achieves a utility score of $0.85$, which we deem adequate to assess the presence of data drifts. On the contrary, an MLP with $5$ layers, $40$ neurons per layer, and a training period of $10$ epochs achieves a utility score of $0.92$ which, despite being higher than $0.85$, does not provide us with better information than the previous configuration.
This is true especially if we consider that the training requirements of more complex MLP are significantly larger compared to models that leverage simpler designs.
VII-B Performance with the ”Fingerprinting” Dataset
@@
Figure 7 depicts the performance, in terms of the score function established by the Eq. 10, for the proposed ALERT method and the baseline techniques described in Section VI-C. The maximum achievable score for this Use Case is $4.5$, which is achieved if all the predictions are correct, and the lowest is $-4.5$ which is obtained if all the predictions are incorrect. We observe that the proposed ALERT method achieves the best score ($4.28$), followed by the CVM ($2.0$) and Anderson Darling ($1.3$) tests. The rest of the methods underperform by a large margin and they obtain negative scores.
TABLE I: Performance in terms of Precision, Recall and F1 of different drift detection methods over the fingerprinting dataset. The proposed ALERT method outperforms all the baseline models, achieving an overall F1-score of $0.9$. In terms of F1-scores, CVM follows with $0.56$, Anderson Darling with $0.41$ and PSI with $0.13$
Methods True positives True negatives False positives False negatives Precision Recall F1
ALERT $9/10$ $19/20$ $1/20$ $1/10$ $0.9$ $0.9$ $0.9$
CVM $8/10$ $10/20$ $10/20$ $2/10$ $0.44$ $0.8$ $0.56$
Anderson Darling $5/10$ $11/20$ $9/20$ $5/10$ $0.35$ $0.5$ $0.41$
PSI $2/10$ $2/20$ $18/20$ $8/10$ $0.1$ $0.2$ $0.13$
Table I depicts the detailed prediction statistics for the four best performing methods on the ”fingerprinting” dataset. The first column of the table represents the methods under examination, while the rest of the columns depict each method’s performance in terms of true positive, true negative, false positive and false negative predictions, as described in Section 10. We have also included $3$ additional columns that present the precision, recall and F1-score of each method. The ALERT method outperforms the baseline techniques since it achieves the best F1-score, namely $0.9$. The CVM follows with $0.56$, the Anderson Darling with $0.41$ and the PSI with $0.13$. Evidently, ALERT manages not only to accurately detect the existence of drift phenomena ($9/10$ true positive score), but also to correctly assess the absence of feature drift ($19/20$ true negative score), thus saving the $M_{0}$ model of unnecessary re-training operations. CVM, which is the best performing baseline method, achieves a bit lower true positive prediction score ($8/10$), but its performance degrades when considering its true negative assessments ($10/20$). As a result, the $M_{0}$ model would be unnecessarily re-trained several times, when using the CVM method to detect feature drifts. The same observation also holds true for both the Anderson Darling and PSI methods.
VII-C Performance with the ”Links” Dataset
Figure 8 illustrates the comparison of the ALERT technique with existing works over the links dataset. The maximum score for each method is $0.48$. The ALERT achieves a score of $0.47$ and clearly outperforms the rest of the techniques, which acquire an equal score of $0.26$, with the exception of Chi-Square that severely under-performs ($0.16$ score). ALERT manages to successfully capture the data distribution properties of the links dataset and to assess with high accuracy the existence of feature drift.
@@ VII-C Performance with the ”Links” Dataset
TABLE II: Performance in terms of Precision, Recall and F1, of different drift detection methods over the Links dataset. ALERT achieves the best F1-score, compared with baseline methods, namely $0.88$. CVM and PSI acquire an F1-score of $0.6$, while Anderson Darling achieves $0.49$.
Methods True positives True negatives False positives False negatives Precision Recall F1
ALERT $4/5$ $/3$ 3 $0/3$ $1/5$ $1.0$ $0.8$ $0.88$
CVM $3/5$ $1/3$ $2/3$ $2/5$ $0.6$ $0.6$ $0.6$
Anderson Darling $2/5$ $2/3$ $1/3$ $3/5$ $0.66$ $0.4$ $0.49$
PSI $3/5$ $1/3$ $2/3$ $2/5$ $0.6$ $0.6$ $0.6$
@@
Table II showcases the performance of top-scoring methods, considering their predictions. Similarly to Table I, the first column depicts the methods under examination which are the ALERT, CVM, Anderson Darling and PSI. The rest of the table columns contain information regarding the true positive, true negative, false positive, false negative, precision, recall and F1 scores of each method. The ALERT technique achieves the best true positive score ($4/5$), followed by CVM and PSI ($3/5$) and Anderson Darling ($2/5$). In terms of true negatives, ALERT comes first with $/3$ 3 correct assessments, Anderson Darling second with $2/3$ and CVM and PSI follow with $1/3$. This has a direct impact to each method’s F1, with ALERT scoring $0.88$, CVM and PSI $0.6$ and Anderson Darling $0.49$. The results we obtain from this evaluation scenario using the Links dataset, are in line with the results we observed under the fingerprinting dataset. The ALERT method not only identifies the existence of drift phenomena, but also it properly assesses their absence as well. This does not necessarily hold true for the baseline methods, which perform worse in their true positive and true negative predictions.
@@
VII-D Ablation Study for the Utility Assessment Component
@@
VII-E Execution time requirements
In Table III we present the execution time requirements of each method under examination, in seconds. The first column of the table refers to the method name, while the next two columns present the time requirements of each technique to perform a feature drift assessment, using the fingerprinting and the links datasets correspondingly. Generally, the methods that exist in the literature are very fast, since they complete their assessments within $0.02$ to $0.4$ seconds. On the other hand, ALERT has larger execution times, ranging from $17.2$ $3.5$ seconds, depending on the use case. This is expected, since ALERT partially trains an MLP model, which severely affects its execution time. Nonetheless, in terms of absolute numbers, we consider the execution time requirements of ALERT affordable for real-world applications, even when large data volumes are involved.
TABLE III: The execution time requirements of the methods under examination for both the fingerprinting and links datasets. The execution time is measured in seconds and is calculated for each method independently. The baseline techniques perform faster than ALERT, which requires $17.2$ seconds to complete on the fingerprinting dataset and $3.5$ seconds on the links dataset.
@@ VII-E Execution time requirements
Although our method is significantly slower compared to other methods, it is not necessarily a significant disadvantage. Gradual drifts usually occur in longer periods of time, from several hours, as noted by [45], or several months, as seen in our fingerprinting dataset. Considering this, our method’s higher execution time can still be regarded as relatively fast, especially in the context of adapting to such evolving conditions.
@@ VIII Conclusion
In this paper we have introduced ALERT—a novel feature drift detection method that comprises representation learning, statistical testing, and utility assessment components. We demonstrated ALERT’s superior performance on two real‑world wireless use cases, fingerprinting and link anomaly detection, where it outperformed ten established methods by ensuring that the AI models keep high F1‑scores, namely of 0.90 and 0.88 even in the presence of feature drift when it is suitably detected and re-training triggered. Beyond raw detection accuracy, our work provides a comprehensive analysis workflow that not only pinpoints when and where feature drift occurs but also quantifies its impact on model performance and informs optimal retraining decisions. By rigorously benchmarking ALERT against standard approaches, we advance the state of the art in feature drift detection for wireless networks and offer practitioners a robust, end‑to‑end solution for maintaining reliable AI models in dynamic radio environments.

### v2

@@ VI-E Performance Metrics
[…]
The function is designed to increase the score for correct feature drift predictions and to decrease for incorrect ones. The larger the feature drift, the bigger score is allocated for correct assessments, and the bigger the penalty is given for incorrect predictions. We assume that each method under examination outputs a prediction on whether the $M_{0}$ should be retrained. We distinguish $4$ different possible scenarios for this assessment:
@@
  • A method’s assessment is false negative and incorrectly identifies the absence of feature drift. The penalty of this error is $F1_{gain}$, which is the $F1$-score that the model would gain if it was retrained with the new data.
@@ VI-F Utility Component Contribution
We perform an analysis of the contribution of each of the following components to the total utility score of the ALERT method: (i) $KS_{(D_{0},D_{1})}$; (ii) $KS_{(R_{0},R_{1})}$; (iii) $PSI_{(D_{0},D_{1})}$; and (iv) $PSI_{(R_{0},R_{1})}$. The aim of this analysis is to validate our initial hypothesis that both the KS and the PSI tests are essential for the calculation of the utility score. Through this, we also aim to analyze the impact of the representation learning technique to the overall utility score of the ALERT. To achieve this, we perform an ablation study by measuring the contributions of each component separately, in terms of percentages ($\%$) for each Use Case, and we present the results we obtain in Section VII-E.
@@ VII Results
TABLE II: The MLP parameters used for the implementation of ALERT.
MLP parameters Values
Layers 3 (Dense, Dense, Dense)
Neurons 20, 20, Number of Classes
Activations Sigmoid, Sigmoid, Softmax
Training epochs 3
Batch size 20
Optimizer Adam
Loss Categorical Cross-Entropy
@@ VII-A Parameter Search for the Representation Learning Components
Figure 6 illustrates the results of the $3$-dimensional parameter exploration, which is conducted for the MLP. We experiment with two datasets in which the feature drift phenomenon is prominent, in order to calibrate the MLP parameters. We should note that the utility function should be represented by a high value, to clearly predict the existence of feature drifts. Results indicate that the increase of model complexity is directly correlated with higher utility scores. This is expected, since more complex models manage to properly capture input data representations and thus, to produce better assessments. Nonetheless, there is an optimal parameter space, over which larger and more complex MLP s do not provide significant contributions to the utility score. Empirically, an MLP with $3$ layers, $20$ neurons per layer, and a training period of $3$ epochs achieves a utility score of $0.85$, which we deem adequate to assess the presence of feature drifts. On the contrary, an MLP with $5$ layers, $40$ neurons per layer, and a training period of $10$ epochs achieves a utility score of $0.92$ which, despite being higher than $0.85$, does not provide us with better information than the previous configuration. This is true especially if we consider that the training requirements of more complex MLP are significantly larger compared to models that leverage simpler designs.
Table II lists the MLP parameters used for the ALERT method. As discussed in the paragraph above, we opted for a small model with $3$ layers, each one of which containing $20$ neurons. Since ALERT is a supervised method, the last layer contains a number of neurons equal to the amount of data classes. We utilise the Sigmoid activation function for the first two layers and the Softmax function for the last layer correspondingly. We train the model for $3$ epochs, using a batch size of $20$, selecting Adam as the optimiser and the Categorical cross-entropy as loss. Due to the small size of the MLP we do not employ any regularization methods, and since the training is conducted for $3$ epochs only, we do not use any early-stopping techniques. In order to avoid temporal leakage, we use the $D_{0}$ dataset for training, and the rest of the datasets for inference. This way, future data points are not accounted for, during the model training.
Figure 7: The behavior of PSI and KS in different drift scenarios (no drift, gradual drift and abrupt drift). PSI excels in identifying abrupt drifts, while KS is good at detecting gradual drifts. The combination of both detectors, along with the proposed representation learning technique give ALERT better coverage for multiple drift scenarios.
VII-B Performance under different feature drift scenarios
Figure 7 illustrates the performance of ALERT under different different drift scenarios. To properly assess the effects of KS and PSI in the design of the utility function in Eq. 7 we opt to visualise their outputs for $3$ scenarios, using the fingerprinting dataset, which is illustrated in Figure 1: (i) when no drift exists in the data; (ii) when a gradual drift occurs; and (iii) when an abrupt drift appears. Evidently, for each scenario, ALERT produces different feature embeddings which are then interpreted by the utility calculation module. As discussed in Section V-C, this module utilises both KS and PSI to assess if a drift phenomenon exists. The results confirm the complementarity of KS and PSI incorporated in Eq. 7. More specifically, the feature analysis conducted by KS achieves good results when gradual drifts are under examination, while in such cases PSI underperforms and thus, it cannot properly identify their presence. The opposite observation for abrupt drifts holds true. There, PSI excels in drift detection, whereas KS often fails to identify abrupt drifts. To solve this, ALERT combines the outputs of both estimators and manages to capture both abrupt and gradual drifts. ALERT, KS and PSI behave similarly when dealing with the Links dataset as well, which is illustrated in Figure 2.
@@ VII-B Performance under different feature drift scenarios
TABLE III: Performance in terms of Precision, FRD, Recall and F1 of different drift detection methods over the fingerprinting dataset. The proposed ALERT method outperforms all the baseline models, achieving an overall F1-score of $0.9$. In terms of F1-scores, ML-AD follows with $0.66$, ML-DDE with $0.64$ and ML-EDC with $0.0.64$
Methods True positives True negatives False positives False negatives Precision FDR Recall F1
ALERT $9/10$ $19/20$ $1/20$ $1/10$ $0.9$ $0.1$ $0.9$ $0.9$
CVM $8/10$ $10/20$ $10/20$ $2/10$ $0.44$ $0.56$ $0.8$ $0.56$
Anderson Darling $5/10$ $11/20$ $9/20$ $5/10$ $0.35$ $0.65$ $0.5$ $0.41$
PSI $2/10$ $2/20$ $18/20$ $8/10$ $0.1$ $0.9$ $0.2$ $0.13$
ML-AD $9/10$ $12/20$ $8/20$ $1/10$ $0.53$ $0.47$ $0.9$ $0.66$
ML-EDC $8/10$ $15/20$ $5/20$ $2/10$ $0.61$ $0.39$ $0.8$ $0.62$
ML-DDE $8/10$ $13/20$ $7/20$ $2/10$ $0.53$ $0.47$ $0.8$ $0.64$
@@
TABLE IV: Performance in terms of Precision, FRD, Recall and F1, of different drift detection methods over the Links dataset. ALERT achieves the best F1-score, namely $0.88$. ML-DDE achieves $0.75$, ML-AD $0.72$, CVM and PSI $0.6$
Methods True positives True negatives False positives False negatives Precision FDR Recall F1
ALERT $4/5$ $3/3$ $0/3$ $1/5$ $1.0$ 0 $0.8$ $0.88$
CVM $3/5$ $1/3$ $2/3$ $2/5$ $0.6$ $0.4$ $0.6$ $0.6$
Anderson Darling $2/5$ $2/3$ $1/3$ $3/5$ $0.66$ $0.34$ $0.4$ $0.49$
PSI $3/5$ $1/3$ $2/3$ $2/5$ $0.6$ $0.4$ $0.6$ $0.6$
ML-AD $4/5$ $1/3$ $2/3$ $1/5$ $0.66$ $0.34$ $0.8$ $0.72$
ML-EDC $2/5$ $3/3$ $0/3$ $3/5$ $1.0$ 0 $0.4$ $0.57$
ML-DDE $3/5$ $2/3$ $0/3$ $2/5$ $1.0$ 0 $0.6$ $0.75$
VII-C Performance with the ”Fingerprinting” Dataset
Figure 8 depicts the performance, in terms of the score function established by the Eq. 10, for the proposed ALERT method and the baseline techniques described in Section VI-C. The maximum achievable score for this Use Case is $4.5$, which is achieved if all the predictions are correct, and the lowest is 4.5-4.5 which is obtained if all the predictions are incorrect. We observe that the proposed ALERT method achieves the best score ($4.28$), followed by the ML-AD ($3.24$), ML-EDC ($2.8$), ML-DDE ($2.6$), CVM ($2.0$) and Anderson Darling ($1.3$) tests. The rest of the methods underperform by a large margin and they obtain negative scores. The superiority of ALERT is due to the learnt representation that keeps only the relevant signals related to drift, and to the design of the utility function in Eq. 7 that provides balanced sensitivity, as discussed in Section V-C. This also contributes to ALERT’s robustness, reducing false positives and false negatives in drift detection.
Table III depicts the detailed prediction statistics for the four best performing methods on the ”fingerprinting” dataset. The first column of the table represents the methods under examination, while the rest of the columns depict each method’s performance in terms of true positive, true negative, false positive and false negative predictions, as described in Section VI-E. We have also included $4$ additional columns that present the precision, false discovery rates (FDR), recall and F1-score of each method. The ALERT method outperforms the baseline techniques since it achieves the best F1-score, namely $0.9$. ML-based methods follow with ML-AD ($0.66$), ML-DDE ($0.64$) and ML-EDC ($0.62$). The rest of the methods perform poorly, as the CVM achieves $0.56$ F1, the Anderson Darling $0.41$ and the PSI $0.13$. Evidently, ALERT manages not only to accurately detect the existence of drift phenomena ($9/10$ true positive score), but also to correctly assess the absence of feature drift ($19/20$ true negative score), thus saving the $M_{0}$ model of unnecessary re-training operations. Also, ALERT minimizes the rate under which it falsely identifies drifts, achieving a low FDR score ($0.1$). ML-AD, which is the best performing baseline method, achieves the same true positive prediction score ($9/10$), but a lower true negative score ($12/20$). Further, ML-AD has a higher FDR score ($0.47$) which would result in unnecessarily re-training operations of the $M_{0}$ model. The same observation also holds true for both the rest of the baseline methods.
VII-D Performance with the ”Links” Dataset
Figure 9 illustrates the comparison of the ALERT technique with existing works over the links dataset. The maximum score for each method is $0.48$. The ALERT achieves a score of $0.47$ and outperforms the rest of the baseline techniques. Distance-based and statistics-based methods acquire an equal score of $0.26$, with the exception of Chi-Square that severely under-performs ($0.16$ score). On the other hand, ML-based techniques perform well, with ML-DDE leading with $0.35$ followed by ML-AD ($0.32$) and ML-EDC ($0.22$). ALERT manages to successfully capture the data distribution properties of the links dataset, containing both gradual and abrupt drifts, and to assess with high accuracy the existence of feature drift.
@@ VII-D Performance with the ”Links” Dataset
Table IV showcases the performance of top-scoring methods, considering their predictions. Similarly to Table III, the first column depicts the methods under examination which are the ALERT, CVM, Anderson Darling, PSI, ML-AD, ML-EDC and ML-DDE. The rest of the table columns contain information regarding the true positive, true negative, false positive, false negative, precision, FDR, recall and F1 scores of each method. The ALERT and the ML-AD technique achieve the best true positive score ($4/5$), followed by ML-DDE, CVM and PSI ($3/5$). In terms of true negatives, ALERT and ML-EDC come first with $3/3$ correct assessments, Anderson Darling and ML-DDE second with $2/3$ and the rest follow with $1/3$. This has a direct impact to each method’s F1, with ALERT scoring $0.88$, closely followed by ML-DDE ($0.75$) and ML-AD ($0.72$), while CVM and PSI score $0.6$. We also observe that all ML-EDC, ML-DDE and ALERT achieve 0 FDR scores and thus, they minimize the unnecessary model re-training operations. Apart from the FDR, the rest of results are in line with the results we observed under the fingerprinting dataset. The ALERT method not only identifies the existence of drift phenomena, but also it properly assesses their absence as well.
@@
VII-E Ablation Study for the Utility Assessment Component
@@
TABLE V: The execution time requirements of the methods under examination for both the fingerprinting and links datasets. The execution time is measured in seconds and is calculated for each method independently.
@@
ML-AD $22$s $4s$
ML-EDC $25$s $5.2s$
ML-DDE $28$s $7s$
VII-F Execution time requirements
In Table V we present the execution time requirements of each method under examination, in seconds. The first column of the table refers to the method name, while the next two columns present the time requirements of each technique to perform a feature drift assessment, using the fingerprinting and the links datasets correspondingly. Generally, the distance-based and statistics-based methods that exist in the literature are very fast, since they complete their assessments within $0.02$ to $0.4$ seconds. On the other hand, ALERT has larger execution times, ranging from $17.2$ - $3.5$ seconds, depending on the use case. This is expected, since ALERT partially trains an MLP model, which severely affects its execution time. Similarly, ML-based methods are slower compared to the others, since during the inference process they utilize a fully trained ML model. We should note that the execution times depicted in Table V correspond to the inference process of the ML-based methods, and they do not account for the model training time requirements. We conclude that, in terms of absolute numbers, the execution time requirements of ALERT are affordable for real-world applications, even when large data volumes are involved.
Further, gradual drifts usually occur in longer periods of time, from several hours, as noted by [48], or several months, as seen in our fingerprinting dataset. Considering this, our method’s higher execution time can still be regarded as relatively fast, especially in the context of adapting to such evolving conditions.
Figure 11: The execution time requirements of the ALERT methods when conducting several drift assessment tasks in parallel.
ALERT’s capacity to scale is illustrated in Figure 11, where we evaluate its performance when executing multiple drift assessments in parallel, simulating large-scale monitoring of diverse data sources. Each dataset contains 16,290 samples and is processed on a single machine with an NVIDIA GeForce GTX 1660 Ti (6 GB GPU) and an Intel i7-10750H (2.6 GHz CPU). The x-axis shows the number of datasets assessed in parallel, while the y-axis indicates the average execution time (in seconds) per assessment. The results show that ALERT scales efficiently, as the difference between 1 and 50 datasets is $\approx$2 seconds of execution overhead.
TABLE VI: The details, advantages and disadvantages of the drift detectors under examination.
Statistical-based methods Distance-based methods ML-based methods ALERT
Method type Unsupervised Supervised Unsupervised Supervised
Parameters p-value threshold Probability bins, cut-off thresholds ML hyperparameters MLP hyperparameters, $T_{s}$
Drift assessment Over data Over data Over features Over features
Low parameter sensitivity
Fast training
Fast inference
Resistant to noise
Do not require regular re-training
High accuracy with small datasets
VII-G Qualitative evaluation of drift detectors
In this section, we discuss the drift detectors which are evaluated throughout this paper. We focus on breaking down their main advantages and disadvantages, and we present their input parameters. Our goal is to analyze the requirements of each technique and to assess their functionality in different scenarios. Under this premise, we also perform an analysis on how ALERT departs from the standard drift detection paradigm and what innovations it brings to the current discourse. Table VI depicts several characteristics of the method groups under examination i.e., statistical-based, distance-based and ML-based.
Both statistical-based and distance-based methods conduct their drift assessments over the input data, are fast to train and support lightweight inference. On the other hand, they require frequent re-training in order to maintain high accuracy, they are not resistant to noise and they underperform when the training dataset is small. Further, they are very sensitive to input parameters which include the p-value threshold (for statistical techniques) and the number of probability bins, as well as the cut-off threshold, over which a drift is detected (for distance-based methods). A key difference between statistical-based and distance-based detectors is that the former are mostly unsupervised methods, while the latter usually fall into the category of supervised methods.
The ML-based techniques, including ALERT, utilise features (instead of raw data) to assess drifts, they have high resistance to noise and they are not very sensitive to their input parameters. On the flip side, their inference process is slower compared to statistical and distance based methods. The ML-based methods that exist in the literature require a significant amount of training time (ranging from minutes to hours), they can work with both labeled and unlabeled data, they need regular re-training to maintain high quality models, and their performance drops when the training dataset size is small. On the contrary, ALERT is a supervised method that does not require frequent re-training operations and works well with both large and smaller datasets.
@@ VIII Conclusion
In this paper we have introduced ALERT—a novel feature drift detection method that comprises representation learning, statistical testing, and utility assessment components. We demonstrated ALERT’s superior performance on two real‑world wireless use cases, fingerprinting and link anomaly detection, where it outperformed ten established methods by ensuring that the AI models keep high F1‑scores, namely of 0.90 and 0.88 even in the presence of feature drift when it is suitably detected and re-training triggered. Beyond raw detection accuracy, our work provides a comprehensive analysis workflow that not only pinpoints when and where feature drift occurs but also quantifies its impact on model performance and informs optimal retraining decisions. By rigorously benchmarking ALERT against standard and state-of-the-art approaches, we advance the state of the art in feature drift detection for wireless networks and offer practitioners a robust, end‑to‑end solution for maintaining reliable AI models in dynamic radio environments.
Although ALERT was evaluated on wireless network data, its architecture is domain-agnostic. Because it operates on learned feature representations rather than on specific radio-signal characteristics, the method can generalize to a wide range of time-series or tabular data domains. This includes applications such as financial transaction monitoring, healthcare sensor analysis, and industrial IoT systems, where data distributions evolve over time. Future work will explore adapting ALERT to these contexts, investigating how representation learning can capture domain-specific dynamics while maintaining robust drift detection across heterogeneous data sources. Additionally, a study of using such detection systems in large scale systems where feature dimensionality or incoming data volumes vary would also be relevant.